If a student completes 7/8 of a math quiz that has forty questions, how many questions did the student answer?
Answer by nerdybill(7384)

If a student completes 7/8 of a math quiz that has forty questions, how many questions did the student answer?
(7/8)(40) = 7*5 = 35 questions